40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3778     Entry time: Mon Oct 25 20:20:59 2010
Author: yuta, Joe 
Type: Update 
Category: CDS 
Subject: DAQ channel number etc 
We split the old SDSEN filters to SDSEN, SDSIDE, SDCOIL last week.
Along with this change, the TP channel number changed unfortunately.
So, we fixed them.
Also, we made FM9 do the output filter analog/digital switching.

What we did:
1. Changed the Simulink logic so that FM9 do the output filter switching, and checked the logic by probing
MAX333A for SDCOILs.

2. After making a new Simulink model and rebuilding, run the following incantation to burt restore filter
settings in /opt/rtcds/caltech/c1/target/c1SYS/c1SYSepics/ (See elog #3706)
sed -i 's/RO \(.*SW[12]R.*\)/\1/' autoBurt.req

3. DAQ channel numbers are listed in C1SYS.ini files in /cvs/cds/rtcds/caltech/c1/chans/daq/.
Channels with # signs are not activated. So, we changed, for example,




for all channels we want to look at.

4. Restarted fb.

- measure TFs and see if input and output filter switchings are working correctly
- make a switch that does all filter switching for all 5 OSEMS or 5 COILS
- put optical lever stuff
- fix offset sliders and offset switch
- put F2A filters in TO_COIL matrix (see elog #3769)
- make a nice graphical screen for MCs (like /cvs/cds/caltech/medm/c1/ioo/C1IOO_ModeCleaner.adl)
- write a script that activates DAQ we need
- make a plan

* SYS=SUS, RMS, MCS, etc
ELOG V3.1.3-